# Learning Spatially Adaptive $\ell_{1}$-Norms Weights for Convolutional Synthesis Regularization

### v2

@@ I Introduction
Contribution. In this work, we propose a synthesis-based spatially adaptive reguralization scheme. We structure our proposal around two main modules. First, we learn adaptive parameter maps that weight the convolutional sparse approximation using pre-learned filters. Then, we solve the corresponding weighted-$\ell_{1}$ problem using the fast iterative shrinkage algorithm (FISTA) [6]. By utilizing pre-computed filters, our formulation avoids the need for alternating minimization while remaining sample-adaptive, as the regularization strength is estimated at inference for each specific sample. Our work is inspired by a recently proposed approach [7] which uses adaptive spatio-temporal Total Variation (TV) regularization maps parametrized by convolutional neural networks (CNNs), which are learned in a supervised fashion through unrolling of the primal-dual hybrid-gradient (PDHG) algorithm [8, 9] ; see also [10] for a recent extension to Total Generalized Variation. Note that employing standard finite-difference discretization filters of the gradient operator corresponds to enforcing a sparsity-based approach with fixed hand-crafted filters within an analysis setting. We report several numerical results for exemplar low-field MRI data, demonstrating quantitative improvements over non-adaptive models and providing quantitative indications (encoded in the estimated maps) about the adequacy of individual filters for signal representation.
@@ II-A Problem Formulation
with $\boldsymbol{\Lambda}\in\mathbb{R}^{NK\times NK}$ being a block-diagonal operator with strictly positive entries, i.e. $\boldsymbol{\Lambda}:=\mathrm{diag}(\boldsymbol{\Lambda}_{1},\ldots,\boldsymbol{\Lambda}_{K})$, so that the term $\|\boldsymbol{\boldsymbol{\Lambda}}\mathbf{s}\|_{1}$ can be written as $\|\boldsymbol{\boldsymbol{\Lambda}}\mathbf{s}\|_{1}:=\sum_{k=1}^{K}\|\mathbf{s}_{k}\|_{\boldsymbol{\Lambda}_{k},1}:=\sum_{k=1}^{K}\sum_{j=1}^{N}\big{|}\boldsymbol{\Lambda}_{k}[j]\cdot\mathbf{s}_{k}[j]\big{|}_{1}$.
@@ II-B Algorithm Unrolling
To learn the parameter maps $\{\boldsymbol{\Lambda}_{k}\}_{k=1}^{K}$, we unroll $T$ iterations of the well-known accelerated proximal gradient descent method FISTA [6]. In particular, we use the version of FISTA presented in [11], which ensures the convergence of the iterates to a minimizer of problem (6). Note that in the FISTA update, the proximal operator of $\tau g(\mathbf{s}):=\tau\|\boldsymbol{\boldsymbol{\Lambda}}\mathbf{s}\|_{1}$ with $\tau>0$ denoting the algorithmic step-size, can be explicitly computed by $\mathrm{prox}_{g}(\mathbf{z})=\mathcal{S}_{\tau\boldsymbol{\Lambda}}(\mathbf{z})$, the componentwise soft-thresholding operator with thresholds defined by the respective values of the corresponding $\boldsymbol{\Lambda}$-map. Furthermore, convergence is guaranteed under the choice $\tau=1/\|\mathbf{B}\|^{2}$.
Figure 1: The proposed pipeline consists of three main blocks: the first one is a high-pass pre-processing to construct the raw data $\mathbf{y}^{\prime}$. The second defines a re-parametrization of adaptive $\{\boldsymbol{\Lambda}_{k}\}_{k}$-maps in terms of a CNN $\text{NET}_{\Theta}$ given the input image $\mathbf{x}_{0}:=\mathbf{A}^{\scriptstyle\boldsymbol{\mathsf{H}}}\mathbf{y}$. The third consists of unrolling $T$ iterations of the FISTA algorithm for solving (8). Note that the pre-processing block includes a parameter $\beta>0$, which is learned together with the network weights $\Theta$.
@@ III-C Training and Implementation Details
The 2D U-Net $\mathbf{u}_{\Theta}$ used in the CNN-block $\mathrm{NET}_{\Theta}$ in (9) has a 2-$K$-$2K$-$4K$-$2K$-$K$ structure with two input-channels (for the real and imaginary parts of the image) and $K$ output channels for $K$ parameter maps $\left\{\mathbf{\Lambda}_{k}\right\}_{k}$. To handle complex-valued data, note that the real and the imaginary parts of the images are approximated using the same filters, i.e. we have $\mathbf{D}\mathbf{s}:=\sum_{k}(d_{k}\ast\mathrm{Re}(\mathbf{s}_{k})+\mathrm{i}\,d_{k}\ast\mathrm{Im}(\mathbf{s}_{k}))$. For training, we used the Adam optimizer [18] with an MSE-loss using a batch-size of 2. For all methods, the scalar regularization parameters ($\lambda$ and $\beta$), which were ”activated” using a Softplus activation function to maintain strict positivity, were assigned a learning rate of $10^{-1}$, while the U-Nets $\mathbf{u}_{\Theta}$, were assigned a learning rate of $10^{-4}$ with a weight decay regularization with weight $10^{-5}$.
@@ IV Results and Discussion
Figure 3 shows an example of images reconstructed for a noise level of $\sigma=0.3$. All methods successfully improved upon the adjoint-based reconstruction, which heavily suffers from noise and blurring artifacts. The methods employing only scalar regularization values $\lambda$ exhibit larger errors at the boundaries and are clearly outperformed by both TV-$\boldsymbol{\Lambda}$ [7] and the proposed CDL-$\boldsymbol{\Lambda}$, which employ learned regularization parameter maps and yield results comparable to the deep learning-based method MoDL [17]. Although similar in terms of SSIM and PSNR, the investigated CDL-$\boldsymbol{\Lambda}$ yields more anatomically accurate images in contrast to TV-$\boldsymbol{\Lambda}$, which, despite the spatially adaptive regularization, still tends to exhibit some piece-wise constant structures. Finally, the investigated CDL-$\boldsymbol{\Lambda}$ exhibits a high visual similarity when compared to MoDL.
Figure 4 summarizes the results obtained on the test set in terms of SSIM and PSNR. The methods involving only scalar regularization parameters (TV-$\lambda$ and CDL-$\lambda$) are outperformed by their corresponding counterparts (TV-$\boldsymbol{\Lambda}$ and CDL-$\boldsymbol{\Lambda}$) involving the learned adaptive spatial regularization, which are competitive even when compared to the deep learning-based MoDL [17].

### v3

@@ I Introduction
Contribution. In this work, we propose a synthesis-based spatially adaptive reguralization scheme. We structure our proposal around two main modules. First, we learn adaptive parameter maps that weight the convolutional sparse approximation using pre-learned filters. Then, we solve the corresponding weighted-$\ell_{1}$ problem using the fast iterative shrinkage algorithm (FISTA) [6]. By utilizing pre-computed filters, our formulation avoids the need for alternating minimization while remaining sample-adaptive, as the regularization strength is estimated at inference for each specific sample. Our work is inspired by a recently proposed approach [7] which uses adaptive spatio-temporal Total Variation (TV) regularization maps parametrized by convolutional neural networks (CNNs), which are learned in a supervised fashion through unrolling of the primal-dual hybrid-gradient (PDHG) algorithm [8, 9] ; see also [10] for a recent extension to Total Generalized Variation. Note that employing standard finite-difference discretization filters of the gradient operator corresponds to enforcing a sparsity-based approach with fixed hand-crafted filters within an analysis setting.
We report several numerical results for exemplar low-field MRI data, demonstrating quantitative improvements over non-adaptive models and providing quantitative indications (encoded in the estimated maps) about the adequacy of individual filters for signal representation.
@@ II-A Problem Formulation
with $\boldsymbol{\Lambda}\in\mathbb{R}^{NK\times NK}$ being a block-diagonal operator with strictly positive entries, i.e. $\boldsymbol{\Lambda}:=\mathrm{diag}(\boldsymbol{\Lambda}_{1},\ldots,\boldsymbol{\Lambda}_{K})$, so that the term $\|\boldsymbol{\boldsymbol{\Lambda}}\mathbf{s}\|_{1}$ can be written as $\|\boldsymbol{\boldsymbol{\Lambda}}\mathbf{s}\|_{1}:=\sum_{k=1}^{K}\|\mathbf{s}_{k}\|_{\boldsymbol{\Lambda}_{k},1}:=\sum_{k=1}^{K}\sum_{j=1}^{N}\big{|}\boldsymbol{\Lambda}_{k}[j]\cdot\mathbf{s}_{k}[j]\big{|}_{1}$. To compute such adaptive $\ell_{1}$ maps, we will make use of an algorithm unrolling strategy. Note that in [11] similar ideas for adaptive least-square regularizers are used.
@@ II-B Algorithm Unrolling
To learn the parameter maps $\{\boldsymbol{\Lambda}_{k}\}_{k=1}^{K}$, we unroll $T$ iterations of the well-known accelerated proximal gradient descent method FISTA [6]. In particular, we use the version of FISTA presented in [12], which ensures the convergence of the iterates to a minimizer of problem (6). Note that in the FISTA update, the proximal operator of $\tau g(\mathbf{s}):=\tau\|\boldsymbol{\boldsymbol{\Lambda}}\mathbf{s}\|_{1}$ with $\tau>0$ denoting the algorithmic step-size, can be explicitly computed by $\mathrm{prox}_{g}(\mathbf{z})=\mathcal{S}_{\tau\boldsymbol{\Lambda}}(\mathbf{z})$, the componentwise soft-thresholding operator with thresholds defined by the respective values of the corresponding $\boldsymbol{\Lambda}$-map. Furthermore, convergence is guaranteed under the choice $\tau=1/\|\mathbf{B}\|^{2}$. As an alternative, iterative rewweighted-$\ell_{1}$ strategies can be used as in [13], with the disadvantage of being less interpretable from the underlying variational perspective.
Figure 1: The proposed pipeline consists of three blocks: the first one is a high-pass pre-processing to construct the raw data $\mathbf{y}^{\prime}$. The second defines a re-parametrization of adaptive $\{\boldsymbol{\Lambda}_{k}\}_{k}$-maps in terms of a CNN $\text{NET}_{\Theta}$ given the input image $\mathbf{x}_{0}:=\mathbf{A}^{\scriptstyle\boldsymbol{\mathsf{H}}}\mathbf{y}$. The third consists of unrolling $T$ iterations of the FISTA algorithm for solving (8). Note that the pre-processing block includes a parameter $\beta>0$, which is learned together with the network weights $\Theta$.
@@ III-C Training and Implementation Details
The 2D U-Net $\mathbf{u}_{\Theta}$ used in the CNN-block $\mathrm{NET}_{\Theta}$ in (9) has a 2-$K$-$2K$-$4K$-$2K$-$K$ structure with two input-channels (for the real and imaginary parts of the image) and $K$ output channels for $K$ parameter maps $\left\{\mathbf{\Lambda}_{k}\right\}_{k}$. To handle complex-valued data, note that the real and the imaginary parts of the images are approximated using the same filters, i.e. we have $\mathbf{D}\mathbf{s}:=\sum_{k}(d_{k}\ast\mathrm{Re}(\mathbf{s}_{k})+\mathrm{i}\,d_{k}\ast\mathrm{Im}(\mathbf{s}_{k}))$. For training, we used the Adam optimizer [20] with an MSE-loss using a batch-size of 2. For all methods, the scalar regularization parameters ($\lambda$ and $\beta$), which were ”activated” using a Softplus activation function to maintain strict positivity, were assigned a learning rate of $10^{-1}$, while the U-Nets $\mathbf{u}_{\Theta}$ were given a learning rate of $10^{-4}$ with a weight decay regularization weight of $10^{-5}$.
@@ IV Results and Discussion
Figure 3 shows an example of images reconstructed for a noise level of $\sigma=0.3$. All methods successfully improved upon the adjoint-based reconstruction, which heavily suffers from noise and blurring artifacts. The methods employing only scalar regularization values $\lambda$ exhibit larger errors at the boundaries and are clearly outperformed by both TV-$\boldsymbol{\Lambda}$ [7] and the proposed CDL-$\boldsymbol{\Lambda}$, which employ learned regularization parameter maps and yield results comparable to the deep learning-based method MoDL [19]. Although similar in terms of SSIM and PSNR, the investigated CDL-$\boldsymbol{\Lambda}$ yields more anatomically accurate images in contrast to TV-$\boldsymbol{\Lambda}$, which, despite the spatially adaptive regularization, tends to exhibit some piece-wise constant structures. Finally, CDL-$\boldsymbol{\Lambda}$ exhibits a high visual similarity when compared to MoDL.
Figure 4 summarizes the results obtained on the test set in terms of SSIM and PSNR. The methods involving only scalar regularization parameters (TV-$\lambda$ and CDL-$\lambda$) are outperformed by both TV-$\boldsymbol{\Lambda}$ and CDL-$\boldsymbol{\Lambda}$ involving the learned adaptive spatial regularization, which are competitive even when compared to the deep learning-based MoDL [19].